OpenStudy (anonymous):

What is the domain of 1/(4 (5-sqrt(x))^(3/2)sqrt(x))

OpenStudy (anonymous):

Im getting 0<x<25 but I dont know how to put that in this format; for example, [0,25)

OpenStudy (mertsj):

\[\frac{1}{4(5-\sqrt{x})^{\frac{3}{2}\sqrt{x}}}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

the sqrt(x) is not an exponent it is just being multiplied

OpenStudy (mertsj):

Well then x can't be negative, 0, 25

OpenStudy (mertsj):

(0,25)U(25,inf)

OpenStudy (anonymous):

okay thanks!

OpenStudy (mertsj):

yw